OpenStudy (anonymous):

solving a impulse, momentum problem. If a ball falls from a roof at 20 degrees and initial speed of 10 m/s what is the velocity of the ball when it hits a table 1 meter below?

OpenStudy (shamim):

U know the x component of the velocity is=10*cos20

OpenStudy (shamim):

Y component of the velicity is=10*sin20

OpenStudy (shamim):

U know x component of ur initial velocity will not change. Because there is no x component of gravitational acceleration g

OpenStudy (shamim):

But y component of initisal velocity will b change

OpenStudy (shamim):

*initial

OpenStudy (shamim):

I think u wanna know the velocity under 1m of initial velocity. Right????

OpenStudy (shamim):

So y component of velocity under 1 m of initial velocity is=Vy

OpenStudy (shamim):

U know Vy^2=u^2-2gh Vy^2=(10*sin20)^2-2*9.8*(-1) Vy=?

OpenStudy (shamim):

Vx=10*cos20

OpenStudy (shamim):

V^2=Vx^2+Vy^2

OpenStudy (shamim):

V=?

OpenStudy (shamim):

Response plz!!!!!!!!

OpenStudy (anonymous):

Here: Vy^2=u^2-2gh Vy^2=(10*sin20)^2-2*9.8*(-1) Vy=? height is -1 but if a consider as going down to be my positive direction I can change it to 1 right?

OpenStudy (anonymous):

so Vx = 9.39 Vy = -7.90 (or 7.90 if we consider going down as the positive direction) V = 12.27 Is this right?

OpenStudy (anonymous):

Now follow up the table is rigid and weights 4kg with a length of 2 meters between the support points a and b and 1 meter from the floor. If the ball has a coefficient of restitution of 0.75. How i can obtain the impulse transferred to the beam and impulse at support points?

OpenStudy (shamim):

Correct!!!!!!!!!!!!!!

OpenStudy (shamim):

Plz go to "ask a question" for any new question posting